# DSO: Aligning 3D Generators with Simulation Feedback for Physical Soundness

### v1

@@ Abstract
Most 3D object generators focus on aesthetic quality, often neglecting physical constraints necessary in applications. One such constraint is that the 3D object should be self-supporting, i.e., remains balanced under gravity. Prior approaches to generating stable 3D objects used differentiable physics simulators to optimize geometry at test-time, which is slow, unstable, and prone to local optima. Inspired by the literature on aligning generative models to external feedback, we propose Direct Simulation Optimization (DSO), a framework to use the feedback from a (non-differentiable) simulator to increase the likelihood that the 3D generator outputs stable 3D objects directly. We construct a dataset of 3D objects labeled with a stability score obtained from the physics simulator. We can then fine-tune the 3D generator using the stability score as the alignment metric, via direct preference optimization (DPO) or direct reward optimization (DRO), a novel objective, which we introduce, to align diffusion models without requiring pairwise preferences. Our experiments show that the fine-tuned feed-forward generator, using either DPO or DRO objective, is much faster and more likely to produce stable objects than test-time optimization. Notably, the DSO framework works even without any ground-truth 3D objects for training, allowing the 3D generator to self-improve by automatically collecting simulation feedback on its own outputs.
@@
Figure 1: Top-left: A state-of-the-art image-to-3D model like TRELLIS often fails to reconstruct 3D objects that can stand under gravity, even when prompted with images of stable objects (e.g., bottom-left). Top-right: Our method, DSO, improves the image-to-3D model via Direct Simulation Optimization, significantly increasing the likelihood that generated 3D objects can stand, both in physical simulation and in real life when 3D printed (bottom-right). The method incurs no additional cost at test time and thus generate such objects in seconds.
@@ 1 Introduction
Given a single image of an object that is stable under gravity, we consider the problem of reconstructing it in 3D. Recent image-to-3D reconstructors [95, 45, 76, 113, 89, 44, 112, 101, 67, 40, 55, 86, 87, 85, 24] have focused on improving the quality of the objects’ 3D geometry and appearance, but not necessarily their physical soundness. As shown in Fig. 2, when prompted with an image of an object which is stable, state-of-the-art generators like TRELLIS [101] and Hunyuan3D 2.0 [87] often fail to produce a stable object. The failure rate is 15% even for objects seen during training and increases significantly for new objects, such as the clock and the motorcycles in Fig. 1.
@@
Stability is a common property of natural and man-made objects and is important in many applications, such as fabrication and simulation [35, 59]. It is, therefore, important to be able to reconstruct 3D objects that satisfy this property.
@@
Works that have explored the generation of physically sound 3D objects [56, 61, 104] have focused on specific object categories, such as furniture. More recent methods like Atlas3D [7] and PhysComp [21] tackle a broader range of object categories. Both methods optimize a 3D model, either from scratch [7] or from the output of an off-the-shelf 3D generator [21], using differentiable physics-based losses that reward stability. To compute these losses, they require differentiable simulators such as [26, 52], which, despite continuous improvements, remain slower and numerically less stable than non-differentiable simulators like [88, 53]. Because of this, Atlas3D and PhysComp are slow and susceptible to local optima and numerical instability.
@@
In this paper, we thus aim to improve a feed-forward 3D generator so that it directly outputs physically stable objects without having to correct them at test time. A naive approach is to use losses similar to those proposed by Atlas3D and PhysComp for feed-forward training instead of test-time optimization, but this would still require a differentiable simulator. Instead, inspired by works on aligning generative models with human preferences [71, 90], we introduce Direct Simulation Optimization (DSO). This simple and effective approach fine-tunes a 3D generator by aligning it with the “preference” provided automatically by an off-the-shelf physics simulator. With this, we explore three research questions: (1) how to use this simulation preference dataset to efficiently fine-tune a 3D generator; (2) how to construct such a dataset without requiring ground-truth 3D data; and (3) whether the fine-tuned generator generalizes well, outputting physically sound 3D objects from image prompts unseen during training.
@@
Our motivation for looking at reward optimization is that stability, like many other physical attributes of an object, is discrete: either an object is stable, or it collapses under gravity. Stability does not distinguish between unstable states regardless of how close they are to becoming stable and is thus difficult to optimize using techniques like gradient descent. In contrast, it is easy to tell if an object is stable or not by using a physics simulator. Hence, we reformulate the problem as a reward-based learning task, where we reward stable outputs and penalize unstable ones. Inspired by direct preference optimization (DPO) [71], we propose an alternative objective, direct reward optimization (DRO), for aligning diffusion models with external preferences. Notably, DRO does not require pairwise preference data for training.
@@
Our second contribution is to show that we can derive the reward signals solely from generated data, which can remove the need to collect many stable 3D objects for training at scale. We do so by generating new 3D assets using the 3D generator itself. These generated 3D assets are then evaluated within a physics simulator, classifying them as stable or unstable. This process allows us to construct a fully automated self-improving pipeline, where the model is trained on its own output assessed by a physics simulator rather than relying on a large dataset of 3D objects.
@@
We show that when integrated with either DPO or DRO as the objective function, our Direct Simulation Optimization framework can steer the output of the 3D generator to align with physical soundness. The final model surpasses previous approaches for physically stable 3D generation on existing evaluation benchmarks [21]. It operates in a feed-forward manner at test time, outperforming heavily engineered solutions like [7, 21] that perform test-time optimization, both in terms of speed and probability of generating a stable object as output. The model also generalizes well to images collected in the wild (Fig. 1). Our experiments show that, in our setting, the proposed DRO objective achieves faster convergence and superior alignment compared to DPO, suggesting that it may be a better candidate for diffusion alignment in general. While our study focuses on stability under gravity, there is nothing special about this particular attribute; the reward-based approach and the self-improving optimization strategy can work, in principle, for any physical attributes that can be assessed via a simulator.
@@ 2 Related Work
Early 3D generators used generative adversarial networks (GANs) [20] and various 3D representations such as point clouds [41, 28], voxel grids [98, 103, 115], view sets [66, 60], NeRF [75, 5, 63, 13, 4], SDF [17], and 3D Gaussian mixtures [97]. However, GANs are difficult to train on a large scale in an ‘open world’ setting. This explains why recent methods have switched to diffusion models [79, 23] while still using the same 3D representations [51, 62, 58, 77, 83, 9], improving training stability and scalability. Others have trained neural networks [30, 106, 100, 107, 99, 29, 39, 27, 84, 6, 8, 82] to directly regress 3D models from 2D images. Authors have also explored scaling 3D reconstruction models [24, 85, 94] on Objaverse [12, 11], improving generalization. DreamFusion [68] and SJC [91] use large-scale image/video generators for 3D generation using score distillation [68, 91, 40, 93, 29, 55, 116]. The works of [45, 76, 36, 47, 54, 22, 113, 96, 48, 18, 86] fine-tune these models for generalizable 3D generation. More recently, authors have introduced latent 3D representations [111, 101, 10] whose distribution can be modeled effectively by denoising diffusion or rectified flow [1, 46, 42]. CLAY [112] and TRELLIS [101] are among the 3D generators trained in this manner, producing superior results compared to methods that piggyback on 2D generation.
These advances have significantly improved the quality of the geometry and appearance of the generated 3D assets, but not necessarily their physical soundness. This limitation reduces their utility in downstream applications like fabrication and simulation. In contrast, we propose a 3D generation approach that explicitly optimizes physical soundness in the guise of stability under gravity.
@@
Early studies have explored methods to predict physical properties from images and videos, like mass [81], shadows [92], materials [109], occlusions [110], and support [78]. While effective in predicting specific physical parameters, these methods do not immediately generalize to 3D reconstruction. Recent works like Physdiff [108], PhysGaussian [102], and PIE-NeRF [15] extended physics-based rendering [26] to NeRF [57] and 3D Gaussian Splatting [32]. These methods focus on modeling the motion of objects instead of their stability under gravity. Like our work, Phys-DeepSDF [56], PhyScene [104], and PhyRecon [61] incorporate explicit physical constraints in 3D reconstruction. However, these methods focus on specific object categories, such as furniture. More related to our work, Atlas3D [7] and PhysComp [21] are not restricted to specific categories; instead, they resort to test-time optimization using carefully designed differentiable, physics-based losses. We solve a similar problem, but in a feed-forward manner using reward-based optimization without requiring to optimize fragile and slow physics-based losses at test time.
@@
The Direct Simulation Optimization (DSO) framework we propose can be trained using direct preference optimization (DPO) [71], a technique originally developed for fine-tuning large language models. Diffusion-DPO [90] first extended DPO to vision diffusion models, enabling direct optimization of human preferences, and was further extended by [16, 43]. While various preference alignment approaches exist [2, 14, 69, 65, 34], DPO has the distinct advantage of not requiring to consult an oracle to compute the reward signal during training, and eschews reward modeling. In this paper, inspired by DPO, we also propose an alternative objective named direct reward optimization (DRO), which does not require pairwise preference data to align the generator.
@@ 3 Method
Given a pre-trained diffusion-based 3D generator $p_{\text{ref}}$ that takes a single image $I$ as input and generates 3D assets $\bm{x}_{0}\sim p_{\text{ref}}(\bm{x}_{0}|I)$, our goal is to learn a new model $p_{\theta}$ that produces more physically sound generations than $p_{\text{ref}}$. We assume access to an oracle $o$ that, given a sample $\bm{x}_{0}$, can output $o(\bm{x}_{0})\in\{0,1\}$, indicating whether $\bm{x}_{0}$ is physically sound. In this paper, we focus on stability under gravity, where $o$ is computed by a physical simulator to determine whether a 3D model $\bm{x}_{0}$ is self-supporting or not.
@@ 3.1 Challenges of Optimizing Physical Soundness
To improve the physical soundness of the generated samples, one approach would be to fine-tune the model with the following objective:
$\max_{\theta}\mathbb{E}$ [${}_{I\sim\mathcal{I},\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}|I)}\left[o(\bm{x}_{0})\right]$
@@
where $\mathcal{I}$ is (the empirical distribution corresponding to) a dataset of image prompts, and $\beta$ is a hyperparameter trading-off the two terms. The first term encourages the generated object $\bm{x}_{0}$ from $p_{\theta}(\bm{x}_{0}|I)$ to be physically sound, while the second term constrains the distribution to remain close to the base model to ensure that the generated geometry remains faithful to the input image $I$.
One challenge in optimizing Sec. 3.1 is that the oracle $o$ is non-differentiable. One approach to address this issue is to reframe the denoising process as a multi-step Markov decision process (MDP) [2] and optimize Sec. 3.1 using reinforcement learning (RL) [73, 74]. However, in our setting, evaluating $o$ is computationally expensive due to the need to run a physical simulation and the overhead introduced by decoding latent 3D representations $\bm{x}_{0}$ into a simulation-ready asset. The decoding process of state-of-the-art 3D generators involves querying dense 3D grid points and extracting a 3D mesh with marching cubes [112, 87], and may even require inference of another geometry generator [101]. These factors make optimization of Sec. 3.1 via RL computationally prohibitive.
@@ 3.2 Formulation as Reward Optimization
We aim to reformulate the objective function to be easier to optimize, specifically eliminating the need to evaluate $o$ during training, while still preserving the intended goals of Sec. 3.1. This is analogous to the goal of text-to-image diffusion model alignment in Diffusion-DPO [90]. In both cases, the reward signal (i.e., evaluation of $o$ by simulation or by collecting human preferences for [90]) is hard to obtain in a scalable way at training time.
@@
$o(\bm{x}_{0})=\beta\mathbb{E}_{p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0},I)}\left[\log\frac{p^{\star}_{\theta}(\bm{x}_{0:T}|I)}{p_{\text{ref}}(\bm{x}_{0:T}|I)}\right]+\beta\log Z(I)$ (2)
for any $I\in\operatorname{supp}(\mathcal{I})$. $Z(I)$ is a normalizing term independent of $p_{\theta}$. The derivation follows [90] and is given in detail in Appendix A.
@@
Given an image dataset $\mathcal{I}$ and 3D models $\mathcal{X}_{I}$ corresponding to each image $I\in\mathcal{I}$, we can then pre-compute $o(\bm{x}_{0})$ for each 3D model $\bm{x}_{0}\in\mathcal{X}_{I}$ to supervise $p_{\theta}$ using Eq. 2, via an $L1$ loss:
$\mathcal{L}\coloneqq$ $(\displaystyle\mathbb{E}_{I\sim\mathcal{I},\bm{x}_{0}\sim\mathcal{X}_{I}}\bigg{[}\bigg{|}o(\bm{x}_{0})-\beta\bigg{(}$3)
@@
To address this issue, we notice that the absolute value of $o(\bm{x}_{0})$ is arbitrary, i.e., we could have used another oracle $o^{\prime}(\bm{x}_{0})\in\{l,u\}$ which evaluates to $l$ for unstable $\bm{x}_{0}$ and $u$ for stable $\bm{x}_{0}$ as long as $l<u$ in Sec. 3.1. In this setting, there exists a choice of $\beta$ that leads to the same optimum $p_{\theta}^{\star}$ as with the original oracle $o$ in Sec. 3.1.
Since we aim to use stochastic gradient descent, which is local and continuous, to optimize $\mathcal{L}$, we may as well choose $l$ and $u$ in a way that within the training compute budget, the sum of $\log Z(I)$ and the expectation over $\bm{x}_{1:T}$ is bounded within $(\frac{l}{\beta},\frac{u}{\beta})$. By doing so, we can remove the absolute value in Eq. 3 and get rid of the terms independent of $p_{\theta}$:
@@
$(1-2o(\bm{x}_{0}))$ $\log\frac{p_{\theta}(\bm{x}_{0:T}|I)}{p_{\text{ref}}(\bm{x}_{0:T}|I)}\bigg{]}$
@@
$\mathcal{L}_{\text{DPO}}\coloneqq$ $-\mathbb{E}_{I\sim\mathcal{I},(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{X}_{I}^{2}}\bigg{[}\log\operatorname{sigmoid}(r(\bm{x}_{0}^{w})-r(\bm{x}_{0}^{l}))\bigg{]}$ (6)
@@
$\mathcal{L}_{\text{DPO}}=$ $-\mathbb{E}_{I\sim\mathcal{I},(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{X}_{I}^{2},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})}$ (7)
$\log\operatorname{sigmoid}\bigg{(}-\beta Tw(t)($
@@
$\quad-\left(\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\theta}(\bm{x}_{t}^{l},t)\|^{2}_{2}-\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\text{ref}}(\bm{x}_{t}^{l},t)\|^{2}_{2}\right)\bigg{)}$

### v2

@@ Abstract
Most 3D object generators prioritize aesthetic quality, often neglecting the physical constraints necessary for practical applications. One such constraint is that a 3D object should be self-supporting, i.e., remain balanced under gravity. Previous approaches to generating stable 3D objects relied on differentiable physics simulators to optimize geometry at test time, which is slow, unstable, and prone to local optima. Inspired by the literature on aligning generative models with external feedback, we propose Direct Simulation Optimization (DSO). This framework leverages feedback from a (non-differentiable) simulator to increase the likelihood that the 3D generator directly outputs stable 3D objects. We construct a dataset of 3D objects labeled with stability scores obtained from the physics simulator. This dataset enables fine-tuning of the 3D generator using the stability score as an alignment metric, via direct preference optimization (DPO) or direct reward optimization (DRO)—a novel objective we introduce to align diffusion models without requiring pairwise preferences. Our experiments demonstrate that the fine-tuned feed-forward generator, using either the DPO or DRO objective, is significantly faster and more likely to produce stable objects than test-time optimization. Notably, the DSO framework functions even without any ground-truth 3D objects for training, allowing the 3D generator to self-improve by automatically collecting simulation feedback on its own outputs.
@@
Figure 1: Top-left: A state-of-the-art image-to-3D model like TRELLIS often fails to reconstruct 3D objects that can stand under gravity even when prompted with images of stable objects (e.g., bottom-left). Top-right: Our method, DSO, improves the image-to-3D model via Direct Simulation Optimization, significantly increasing the likelihood that generated 3D objects can stand, in physical simulation and in real-life, when 3D printed (bottom-right). The method incurs no additional cost at test time, and can thus generate such objects in seconds.
@@ 1 Introduction
Given a single image of an object that is stable under gravity, we consider the problem of reconstructing it in 3D. Recent image-to-3D reconstructors [95, 45, 76, 113, 89, 44, 112, 101, 67, 40, 55, 86, 87, 85, 24] have focused on improving the quality of objects’ 3D geometry and appearance, but not necessarily their physical soundness. As shown in Fig. 2, when prompted with an image of a stable object, state-of-the-art generators like TRELLIS [101] and Hunyuan3D 2.0 [87] often fail to produce a stable object in 3D. The failure rate is 15% even for objects seen during training and increases significantly for new objects, such as the clock and motorcycles in Fig. 1.
@@
Stability is a common property of natural and man-made objects and is important in many applications, such as fabrication and simulation [35, 59]. It is, therefore, important to reconstruct 3D objects that satisfy this property.
@@
Previous works on generating physically sound 3D objects [56, 61, 104] have focused on specific object categories, such as furniture. More recent methods like Atlas3D [7] and PhysComp [21] tackle a broader range of object categories. Both methods optimize a 3D model, either from scratch [7] or from the output of an off-the-shelf 3D generator [21], using differentiable physics-based losses that reward stability. To compute these losses, they require differentiable simulators such as [26, 52], which, despite continuous improvements, remain slower and numerically less stable than non-differentiable simulators like [88, 53]. As a result, Atlas3D and PhysComp are slow and susceptible to local optima and numerical instability.
@@
In this paper, we aim to improve a feed-forward 3D generator so that it directly outputs physically stable objects without requiring test-time corrections. A naïve approach would be to use losses similar to those proposed by Atlas3D and PhysComp for feed-forward training instead of test-time optimization, but this would still require a differentiable simulator. Instead, inspired by works on aligning generative models with human preferences [71, 90], we introduce Direct Simulation Optimization (DSO). This simple and effective approach fine-tunes a 3D generator by aligning it with the “preference” provided automatically by an off-the-shelf physics simulator. With this, we explore three research questions: (1) How to use this simulation preference dataset to fine-tune a 3D generator efficiently; (2) How to construct such a dataset without requiring ground-truth 3D data; and (3) Whether the fine-tuned generator generalizes well, outputting physically sound 3D objects from image prompts unseen during training.
@@
Our motivation for using reward optimization is that stability, like many other physical attributes of an object, is discrete: either an object is stable, or it collapses under gravity. Stability does not distinguish between unstable states regardless of how close they are to becoming stable, making it difficult to optimize using techniques like gradient descent. In contrast, it is easy to determine whether an object is stable or not using a physics simulator. Hence, we reformulate the problem as a reward-based learning task, where we reward stable outputs and penalize unstable ones. Inspired by direct preference optimization (DPO) [71], we propose an alternative objective, direct reward optimization (DRO), for aligning diffusion models with external preferences. Notably, DRO does not require pairwise preference data for training.
@@
Our second contribution is to show that we can derive reward signals solely from generated data, eliminating the need to collect large datasets of stable 3D objects for training at scale. We achieve this by generating new 3D assets using the 3D generator itself. These generated 3D assets are then evaluated within a physics simulator, classifying them as stable or unstable. This process allows us to construct a fully automated self-improving pipeline, where the model is trained on its own output, assessed by a physics simulator rather than relying on a large dataset of 3D objects.
@@
We show that, when integrated with either DPO or DRO as the objective function, our Direct Simulation Optimization framework can steer the output of the 3D generator to align with physical soundness. The final model surpasses previous approaches for physically stable 3D generation on existing evaluation benchmarks [21]. It operates in a feed-forward manner at test time, outperforming heavily engineered solutions like [7, 21] that perform test-time optimization, both in terms of speed and probability of generating a stable object as output. The model also generalizes well to images collected in the wild (Fig. 1).
Our experiments show that, in our setting, the proposed DRO objective achieves faster convergence and superior alignment compared to DPO, suggesting that it may be a better candidate for diffusion alignment in general. While our study focuses on stability under gravity, the reward-based approach and the self-improving optimization strategy can, in principle, be applied to any physical attributes that can be assessed via a simulator.
@@ 2 Related Work
Early 3D generators used generative adversarial networks (GANs) [20] and various 3D representations such as point clouds [41, 28], voxel grids [98, 103, 115], view sets [66, 60], NeRF [75, 5, 63, 13, 4], SDF [17], and 3D Gaussian mixtures [97]. However, GANs are challenging to train on a large scale in an ‘open world’ setting. This explains why recent methods have shifted to diffusion models [79, 23], which use the same 3D representations [51, 62, 58, 77, 83, 9] while improving training stability and scalability. Other approaches train neural networks [30, 106, 100, 107, 99, 29, 39, 27, 84, 6, 8, 82] to directly regress 3D models from 2D images. Researchers have also explored scaling 3D reconstruction models [24, 85, 94] on Objaverse [12, 11], improving generalization. DreamFusion [68] and SJC [91] leverage large-scale image/video generators for 3D generation using score distillation [68, 91, 40, 93, 29, 55, 116]. The works of [45, 76, 36, 47, 54, 22, 113, 96, 48, 18, 86] fine-tune these models for generalizable 3D generation. More recently, researchers have introduced latent 3D representations [111, 101, 10] whose distributions can be effectively modeled by denoising diffusion or rectified flow [1, 46, 42]. CLAY [112] and TRELLIS [101] are among the 3D generators trained in this manner, producing superior results compared to methods that rely on 2D generation.
These advances have significantly improved the quality of the geometry and appearance of generated 3D assets, but not necessarily their physical soundness. This limitation reduces their utility in downstream applications like fabrication and simulation. In contrast, we propose a 3D generation approach that explicitly optimizes physical soundness, specifically stability under gravity.
@@
Early studies explored methods to predict physical properties from images and videos, such as mass [81], shadows [92], materials [109], occlusions [110], and support [78]. While effective in predicting specific physical parameters, these methods do not generalize directly to 3D reconstruction. Recent works like Physdiff [108], PhysGaussian [102], and PIE-NeRF [15] extend physics-based rendering [26] to NeRF [57] and 3D Gaussian Splatting [32]. These methods focus on modeling the motion of objects rather than their stability under gravity. Similar to our work, Phys-DeepSDF [56], PhyScene [104], and PhyRecon [61] incorporate explicit physical constraints in 3D reconstruction. However, these methods are limited to specific object categories, such as furniture. More related to our work, Atlas3D [7] and PhysComp [21] are not restricted to specific categories; instead, they rely on test-time optimization using carefully designed differentiable, physics-based losses. We address a similar problem but in a feed-forward manner using reward-based optimization, avoiding the need for fragile and slow physics-based losses at test time.
@@
The Direct Simulation Optimization (DSO) framework we propose can be trained using Direct Preference Optimization (DPO) [71], a technique initially developed for fine-tuning large language models. Diffusion-DPO [90] first extended DPO to vision diffusion models, enabling direct optimization of human preferences, and was further extended by [16, 43]. While various preference alignment approaches exist [2, 14, 69, 65, 34], DPO has the distinct advantage of not requiring an oracle to compute the reward signal during training and avoids the need for reward modeling. Inspired by DPO, we also propose an alternative objective named direct reward optimization (DRO), which does not require pairwise preference data to align the generator.
@@ 3 Method
Given a pre-trained diffusion-based 3D generator $p_{\text{ref}}$ that takes a single image $I$ as input and generates 3D assets $\bm{x}_{0}\sim p_{\text{ref}}(\bm{x}_{0}|I)$, our goal is to learn a new model $p_{\theta}$ that produces more physically sound generations than $p_{\text{ref}}$. We assume access to an oracle $o$ that, given a sample $\bm{x}_{0}$, outputs $o(\bm{x}_{0})\in\{0,1\}$, indicating whether $\bm{x}_{0}$ is physically sound. In this paper, we focus on stability under gravity, where $o$ is computed by a physics simulator to determine whether a 3D model $\bm{x}_{0}$ is self-supporting.
@@ 3.1 Challenges of Optimizing Physical Soundness
To improve the physical soundness of the generated samples, one approach is to fine-tune the model with the following objective:
$\max_{\theta}\mathbb{E}$ $_{I\sim\mathcal{I},\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}|I)}\left[o(\bm{x}_{0})\right$]
@@
where $\mathcal{I}$ is the empirical distribution of a dataset of image prompts, and $\beta$ is a hyperparameter trading off the two terms. The first term encourages the generated object $\bm{x}_{0}$ from $p_{\theta}(\bm{x}_{0}|I)$ to be physically sound, while the second term constrains the distribution to remain close to the base model to ensure that the generated geometry remains faithful to the input image $I$.
A key challenge in optimizing Sec. 3.1 is that the oracle $o$ is non-differentiable. One approach to address this issue is to reframe the denoising process as a multi-step Markov decision process (MDP) [2] and optimize Sec. 3.1 using reinforcement learning (RL) [73, 74]. However, in our setting, evaluating $o$ is computationally expensive due to the need to run a physical simulation and the overhead introduced by decoding latent 3D representations $\bm{x}_{0}$ into simulation-ready assets. The decoding process of state-of-the-art 3D generators involves querying dense 3D grid points and extracting a 3D mesh with marching cubes [112, 87], and may even require inference of another geometry generator [101]. These factors make optimization of Sec. 3.1 via RL computationally prohibitive.
@@ 3.2 Formulation as Reward Optimization
We aim to reformulate the objective function to be easier to optimize, specifically eliminating the need to evaluate $o$ during training, while still preserving the intended goals of Sec. 3.1. This is analogous to the goal of text-to-image diffusion model alignment in Diffusion-DPO [90]. In both cases, the reward signal (i.e., evaluation of $o$ by simulation or by collecting human preferences for [90]) is hard to obtain in a scalable way during training.
@@
$o(\bm{x}_{0})=\beta\mathbb{E}_{p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0},I)}\left[\log\frac{p^{\star}_{\theta}(\bm{x}_{0:T}|I)}{p_{\text{ref}}(\bm{x}_{0:T}|I)}\right]+\beta\log Z(I),$ (2)
for any $I\in\operatorname{supp}(\mathcal{I})$, where $Z(I)$ is a normalizing term independent of $p_{\theta}$. The derivation follows [90] and is detailed in Appendix A.
@@
Given an image dataset $\mathcal{I}$ and 3D models $\mathcal{X}_{I}$ corresponding to each image $I\in\mathcal{I}$, we can pre-compute $o(\bm{x}_{0})$ for each 3D model $\bm{x}_{0}\in\mathcal{X}_{I}$ to supervise $p_{\theta}$ using Eq. 2, via an $L1$ loss:
$\mathcal{L}\coloneqq$ $\mathbb{E}_{I\sim\mathcal{I},\bm{x}_{0}\sim\mathcal{X}_{I}}\bigg{[}\bigg{|}o(\bm{x}_{0})-\beta\bigg{(}$ (3)
@@
To address this issue, we notice that the absolute value of $o(\bm{x}_{0})$ is arbitrary, i.e., we could use another oracle $o^{\prime}(\bm{x}_{0})\in\{l,u\}$ which evaluates to $l$ for unstable $\bm{x}_{0}$ and $u$ for stable $\bm{x}_{0}$, as long as $l<u$ in Sec. 3.1. In this setting, there exists a choice of $\beta$ that leads to the same optimum $p_{\theta}^{\star}$ as with the original oracle $o$ in Sec. 3.1.
Since we aim to use stochastic gradient descent, which is local and continuous, to optimize $\mathcal{L}$, we may as well choose $l$ and $u$ such that, within the training compute budget, the sum of $\log Z(I)$ and the expectation over $\bm{x}_{1:T}$ is bounded within $(\frac{l}{\beta},\frac{u}{\beta})$. By doing so, we can remove the absolute value in Eq. 3 and get rid of the terms independent of $p_{\theta}$:
@@
$(1-2o(\bm{x}_{0}))$ $\log\frac{p_{\theta}(\bm{x}_{0:T}|I)}{p_{\text{ref}}(\bm{x}_{0:T}|I)}\bigg{]}.$
@@
$\mathcal{L}_{\text{DPO}}\coloneqq$ $-\mathbb{E}_{I\sim\mathcal{I},(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{X}_{I}^{2}}\bigg{[}\log\operatorname{sigmoid}(r(\bm{x}_{0}^{w})-r(\bm{x}_{0}^{l}))\bigg{]},$ (6)
@@
$\mathcal{L}_{\text{DPO}}={}$ $-\mathbb{E}_{\begin{subarray}{c}I\sim\mathcal{I},(\bm{x}_{0}^{w},\bm{x}_{0}^{l})\sim\mathcal{X}_{I}^{2},t\sim\mathcal{U}(0,T),\bm{x}_{t}^{w}\sim q(\bm{x}_{t}^{w}|\bm{x}_{0}^{w}),\bm{x}_{t}^{l}\sim q(\bm{x}_{t}^{l}|\bm{x}_{0}^{l})\end{subarray}}$ (7)
$\log\operatorname{sigmoid}\bigg{(}-\beta Tw(t)\Big{(}$
@@
$\quad-\left(\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\theta}(\bm{x}_{t}^{l},t)\|^{2}_{2}-\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\text{ref}}(\bm{x}_{t}^{l},t)\|^{2}_{2}\right)\Big{)}\bigg{)},$